OpenStudy (anonymous):

I new to this.... But the thing asks for me to find the improper fraction of 5 and 3/10.... It sounds easy, and probably is, but I'm stump. O.o

OpenStudy (unklerhaukus):

convert 5 into tenths

OpenStudy (anonymous):

50? :L

OpenStudy (ahsome):

Equation: \[5\frac{3}{10}\] The end goal is to remove the 5 and make the numerator bigger than the denominator. First, multiply the whole number and the denominator: \[5*10\] \[50\] Add this number to the numerator, while removing the whole number: \[5\frac{3}{10}\] \[\frac{50+3}{10}\] \[\frac{53}{10}\] TADA

OpenStudy (anonymous):

O_O

OpenStudy (ahsome):

Read the instructions. If anything doesn't make sense, just say so.

OpenStudy (unklerhaukus):

yeah 5 is 50 tenths, so 5 + 3/10 = 50/10 + 3/10 = (50+3)/10 =

OpenStudy (anonymous):

Ugh, I probably look pretty stupid right now....

OpenStudy (ahsome):

No, its fine. Just say where you don't understand

OpenStudy (ahsome):

Ok, you know where they say 5? That just means there are 5 other \(\frac{10}{10}\), they just left it out. So,, instead of saying: \[5\frac{3}{10}\] We can say: \[\frac{10}{10}+\frac{10}{10}+\frac{10}{10}+\frac{10}{10}+\frac{10}{10}+\frac{3}{10}\]

OpenStudy (anonymous):

Oh, I guess I get it now... Thanks

OpenStudy (ahsome):

Now, we can just add up all the numerators: \[\frac{10+10+10+10+10+3}{10}\] \[\frac{53}{10}\]

OpenStudy (ahsome):

No problem, @Rhi. :D

OpenStudy (ahsome):

If you think either me or @UnkleRhaukus, please press the best response button on the right. Thanks :)

OpenStudy (ahsome):

If you still don't understand @Rhi, just say so